OpenStudy (anonymous):

which integer added to -6 gives a sum of 8?

OpenStudy (anonymous):

A= the integer A+(-6)=8 A-6+6=8+6 A=14

OpenStudy (anonymous):

k thanx alot

OpenStudy (anonymous):

NP

OpenStudy (anonymous):

thanx bye

OpenStudy (anonymous):

wait y do u add 6+6?

OpenStudy (anonymous):

A-6=8 Add 6 to both side of the equation A-6+6=8+6

OpenStudy (anonymous):

oh now i get it better thanx

OpenStudy (anonymous):

k thanx bye